A generalization of Eneström-Kakeya theorem

V. K. Jain, Kharagpur, India

Keywords: zeros, infinite series with complex coefficients AMS Subject classification: 30C15, 30C10

Abstract: A generalization of Eneström-Kakeya theorem (well known in the theory of the distribution of zeros of polynomials) in relation to infinite series with complex coefficients has been obtained. The result is best possible.